# Entanglement Entropy of Free Fermions in Timelike Slices

### v3

@@ II.1 Definition in the Heisenberg picture
We now show how sub-Hilbert space $h_{A}$ can be identified for free fermion models. Consider a lattice model (in any dimension) with a free fermion Hamiltonian of the fermion bilinear form
$H=\sum_{i,j}c^{\dagger}_{\mathbf{r}_{i}}h_{ij}c_{\mathbf{r}_{j}}\ .$ (4)
Assume each site $\mathbf{r}_{m}$ has one fermion degree of freedom with annihilation operator $c_{\mathbf{r}_{m}}$ and creation operator $c_{\mathbf{r}_{m}}^{\dagger}$ at time $t=0$, with anti-commutators $\{c_{\mathbf{r}_{m}},c_{\mathbf{r}_{n}}^{\dagger}\}=\delta_{mn}$. In an arbitrary slice $A$ with $K$ discrete spacetime points $(\mathbf{r}_{n},t_{n})$, each annihilation operator $c_{\mathbf{r}_{n}}(t_{n})=e^{iHt_{n}}c_{\mathbf{r}_{n}}e^{-iHt_{n}}$ at point $(\mathbf{r}_{n},t_{n})$ is the linear combination of $c_{\mathbf{r}_{m}}$ at time $t=0$. Accordingly, the anti-commutations between $c_{\mathbf{r}_{n}}(t_{n})$ and $c_{\mathbf{r}_{m}}^{\dagger}(t_{m})$ are not delta functions, but are still numbers:
@@
$B_{mn}$ gives a $K\times K$ non-negative Hermitian commutation matrix $B$ for the $K$ points in slice $A$. By rewriting $B=Q_{1}\Lambda Q_{1}^{\dagger}$ with $\Lambda$ diagonal and $Q_{1}$ unitary, we can define a matrix $M=Q_{2}\Lambda^{-1/2}Q_{1}^{\dagger}$ where $Q_{2}$ can be an arbitrary unitary matrix, such that $M$ satisfies $M^{\dagger}M=B^{-1}$. Particularly, by QL decomposition, we choose $Q_{2}$ such that $M$ is a lower triangular matrix for later purposes of path integral formalism. We can then write down an orthonormal fermion basis:
$d_{m}=\sum_{n\in A}M_{mn}c_{\mathbf{r}_{n}}(t_{n})\ ,$ (6)
@@
and thus form an orthonormal fermion basis. $h_{A}$ is then the Hilbert space of the $K$ fermion degrees of freedom $d_{m}$, which has a dimension $\text{dim}[h_{A}]=2^{K}$. Any correlations among $c_{\mathbf{r}_{n}}(t_{n})$ and $c_{\mathbf{r}_{m}}^{\dagger}(t_{m})$ are completely determined within $h_{A}$. We note that our method for identifying sub-Hilbert space $h_{A}$ here is similar to the bulk Hilbert space reconstruction most recently studied in Ref. [43].
As an illustrative example, consider a lattice of 3 sites with corresponding fermion annihilation operators $c_{1},c_{2},c_{3}$, respectively, and assume spacetime slice $A$ contains two spacetime points: $(x_{1},t_{1})=(x_{1},0)$ and $(x_{2},t_{2})$, with $t_{1}=0$ and $t_{2}\neq 0$. Assume the free Hamiltonian $H$ and time $t_{2}$ are such that, the fermion operators at these two points are
@@
When $\rho_{\text{tot}}$ is a mixed state, it is known that the entanglement entropy $S_{A}$ of sub-Hilbert space $h_{A}$ (Eq. 15) is not a measure of the entanglement between $h_{A}$ and $h_{A^{c}}$, but is only a von-Neumann entropy. Instead, the entanglement between sub-Hilbert space $h_{A}$ of slice $A$ and its complement $h_{{A^{c}}}$ can be characterized by the mutual information:
@@
where $S_{A^{c}}=-\text{tr}(\rho_{A^{c}}\ln\rho_{A^{c}})$ is the entanglement entropy of subsystem $A^{c}$, and $S_{\text{tot}}=-\text{tr}(\rho_{\text{tot}}\ln\rho_{\text{tot}})$ is the entanglement entropy of the entire system. For a pure state $\rho_{\text{tot}}$, one has $S_{\text{tot}}=0$ and $\mathcal{I}=2S_{A}$.
@@ II.2 Path integral formalism
To provide an understanding of the above definition of entanglement entropy from a spacetime local perspective, in this subsection we rewrite it in terms of coherent state path integral formalism.
@@
II.3 Model and definitions
We note that our definition of entanglement entropy in generic spacetime slice above holds for any free fermion models in any spacetime dimensions. For simplicity, hereafter, we investigate the spacetime slice entanglement entropy for the 1D tight-binding free fermion model with $L$ sites and periodic boundary condition:
@@ II.3 Model and definitions
III TIME-DIRECTION SLICE
We now consider the case of slice $A$ being a time-direction line segment of total time $t$ on a fixed site $\mathbf{r}_{n}=0$, containing $K$ equally spaced points $(\mathbf{r}_{n},t_{n})=(0,(n-1)\tau)$, where $1\leq n\leq K$, and $\tau=\frac{t}{K-1}$ is the spacing between adjacent times, as shown in Fig. 1b.
We first study the entanglement entropy $S_{A}$ of thermal states $\rho_{\text{tot}}$ at temperature $T$ and filling $\nu$ in slice $A$, and then study that for non-eigenstates. We set $L\gg K$, for which $S_{A}$ is independent of $L$.
III.1 Zero temperature
At temperature $T=0$, $\rho_{\text{tot}}$ is a pure state (ground state with a Fermi sea). For fixed total time $t$, as shown in Fig. 3a, within the range $K<t/\tau_{0}$, or equivalently $\tau>\tau_{0}$, $S_{A}$ shows a triangular (trapezoid) shape curve with respect to $K$ for filling $\nu=0.5$ ($\nu\neq 0.5$); while in the range $K>t/\tau_{0}$, or $\tau<\tau_{0}$, $S_{A}$ approaches a stabilized value. We further examine the total time $t$ dependence of the stabilized value of $S_{A}$ by fixing the time separation $\tau$ and changing the number of points $K=\frac{t}{\tau}+1$. Interestingly, we find $S_{A}$ is approximately linear in $\ln t$, where the slope $\left(\frac{\partial S_{A}}{\partial\ln K}\right)_{\tau}\approx\left(\frac{\partial S_{A}}{\partial\ln t}\right)_{\tau}$ is $\frac{1}{3}$ at $\tau=\tau_{0}$ [51], and quickly saturates to $\frac{1}{6}$ when $\tau<\tau_{0}$ (Fig. 3b). In addition, $S_{A}$ shows an oscillation of period $\text{max}(\frac{1}{\nu},\frac{1}{1-\nu})$ with respect to $K$ [51] at filling $\nu$. These conclusions are insensitive to the energy dispersion [51].
@@ III.1 Zero temperature
Figure 3: (a)-(c) Entanglement entropy $S_{A}$ of model Eq. 32 with $L=500$ in time-direction slice of length $t$ with $K$ points separated by $\tau=t/(K-1)$, for (a) different filling $\nu$ given in the legend at temperature $T=0$ and fixed $1682\pi t/\tau_{0}=$ 168; (b) $S_{A}$ versus $\ln t$ for different fillings $\nu$ (see legend) at $T=0$ and fixed $\tau=\frac{2}{\pi}\tau_{0}$; (c) finite temperature $S_{A}$ with fixed $1682\pi t/\tau_{0}=$ 168, at fillings $\nu=0.5$ (solid) and $\nu=0.3$ (dashed). For each $\nu$, the four curves from low to high have $T/E_{0}=0,0.075,0.15,\infty$, respectively. (d) The ancillary dispersion for explaining $S_{A}$ when $\tau<\tau_{0}$ (“anc” for ancillary, “emp” for empty). (e)-(f) Mutual information $\mathcal{I}$ with $L=200$ at $\nu=0.5$ and temperatures $T/E_{0}$ given in the legends while fixing (e) $1682\pi t/\tau_{0}=$ 168 and (f) $\tau=\frac{5}{2\pi}\tau_{0}$, respectively.
@@
where $n_{F}(E)$ is the Fermi-Dirac distribution function at filling $\nu$.
(i) When $\tau\gg\tau_{0}$, or $K\ll t/\tau_{0}$, one has
@@
where the off-diagonal elements are suppressed by their fast oscillating integrands. This yields $D=M^{*}CM^{T}\approx\nu I$, and by Eq. 15 one has
@@
This fits well with the slope $\left(\frac{\partial S_{A}}{\partial K}\right)_{t}$ at small $K$ in Fig. 3a [51]. Moreover, if one approximates $\Omega(E)$ as uniform:
$\Omega(E)=E_{0}^{-1}\Theta\left(E_{0}/2-|E|\right)\ ,$ (38)
where $\Theta(x)=1$ ($=0$) when $x>0$ ($x<0$), Eq. 37 becomes exact when $\nu=\frac{p}{q}$ ($p,q\in\mathbb{N}^{+}$ coprime) and $\tau=jq\tau_{0}$ ($j\in\mathbb{N}^{+}$). Eq. 37 also implies that $S_{A}\propto t=(K-1)\tau$ for fixed $\tau\gg\tau_{0}$. $S_{A}$ in this limit is similar to the tensor network temporal entanglement entropy in [49, 50], although the definitions differ [51].
(ii) When $\tau=\tau_{0}$, or $K=\frac{t}{\tau_{0}}+1$, taking the uniform $\Omega(E)$ approximation in Eq. 38, one has exactly

### v4

@@ II.1 Definition in the Heisenberg picture
We now show how sub-Hilbert space $h_{A}$ can be identified for free fermion models. Consider a lattice model (in any dimension) where each site $\mathbf{r}_{m}$ has one fermion degree of freedom with annihilation operator $c_{\mathbf{r}_{m}}$ and creation operator $c_{\mathbf{r}_{m}}^{\dagger}$ at time $t=0$, with anti-commutators $\{c_{\mathbf{r}_{m}},c_{\mathbf{r}_{n}}^{\dagger}\}=\delta_{mn}$. Assume the model has a free fermion Hamiltonian of the fermion bilinear form
$H=\sum_{i,j}c^{\dagger}_{\mathbf{r}_{i}}h_{ij}c_{\mathbf{r}_{j}}\ ,$ (4)
which has charge conservation (the case without charge conservation will be discussed in Sec. II.3). In an arbitrary slice $A$ with $K$ discrete spacetime points $(\mathbf{r}_{n},t_{n})$, each annihilation operator $c_{\mathbf{r}_{n}}(t_{n})=e^{iHt_{n}}c_{\mathbf{r}_{n}}e^{-iHt_{n}}{=\sum_{j}\xi_{nj}c_{\mathbf{r}_{j}}}$ at point $(\mathbf{r}_{n},t_{n})$ is the linear combination of $c_{\mathbf{r}_{m}}$ at time $t=0$. Accordingly, the anti-commutations between $c_{\mathbf{r}_{n}}(t_{n})$ and $c_{\mathbf{r}_{m}}^{\dagger}(t_{m})$ are not delta functions, but are still numbers:
@@
$B_{mn}$ gives a $K\times K$ non-negative Hermitian commutation matrix $B$ for the $K$ points in slice $A$. By rewriting $B=Q_{1}\Lambda Q_{1}^{\dagger}$ with $\Lambda$ diagonal and $Q_{1}$ unitary, we can define a matrix $M=Q_{2}\Lambda^{-1/2}Q_{1}^{\dagger}$ where $Q_{2}$ can be an arbitrary unitary matrix, such that $M$ satisfies
$M^{\dagger}M=B^{-1}.$ (6)
Particularly, by QL decomposition, we choose $Q_{2}$ such that $M$ is a lower triangular matrix for later purposes of path integral formalism. We can then write down an orthonormal fermion basis (in terms of the zero-time on-site basis $c_{\mathbf{r}_{j}}$):
$d_{m}=\sum_{n\in A}M_{mn}c_{\mathbf{r}_{n}}(t_{n})=\sum_{n\in A}\sum_{j}M_{mn}\xi_{nj}c_{\mathbf{r}_{j}}\ ,$ (7)
@@
and thus form an orthonormal fermion basis. $h_{A}$ is then the Hilbert space of the $K$ fermion degrees of freedom $d_{m}$, which has a dimension $\text{dim}[h_{A}]=2^{K}$. More explicitly, the basis of the sub-Hilbert space $h_{A}$ is given by the $2^{K}$ Fock states $\prod_{m}(d_{m}^{\dagger})^{\zeta_{m}}|0\rangle$ for $\zeta_{m}\in\{0,1\}$, $1\leq m\leq K$. In contrast, the total Hilbert space of the system $h_{\text{tot}}$ has a basis given by the $2^{L}$ Fock states $\prod_{m}(c_{\mathbf{r}_{n}}^{\dagger})^{\xi_{n}}|0\rangle$ for $\xi_{n}\in\{0,1\}$, where $L$ is the total number of sites of the system. This allows one to explicitly decompose the total Hilbert space into the subsystem of slice $A$, and its complement $A^{c}$ spanned by the Hilbert space of fermion operators orthogonal to $d_{m}$ in Eq. 7:
$h_{\text{tot}}=h_{A}\otimes h_{A^{c}}\ .$ (9)
Any correlations among $c_{\mathbf{r}_{n}}(t_{n})$ and $c_{\mathbf{r}_{m}}^{\dagger}(t_{m})$ are completely determined within sub-Hilbert space $h_{A}$. We note that our method for identifying sub-Hilbert space $h_{A}$ here is similar to the bulk Hilbert space reconstruction most recently studied in Ref. [43].
As an illustrative example, consider a lattice of 3 sites with corresponding fermion annihilation operators $c_{1},c_{2},c_{3}$, respectively, and assume spacetime slice $A$ contains two spacetime points: $(x_{1},t_{1})=(x_{1},0)$ and $(x_{2},t_{2})$, with $t_{1}=0$ and $t_{2}\neq 0$. For concreteness, assume the free Hamiltonian is $H=\sum_{m,n}c_{m}^{\dagger}\mathcal{H}_{mn}c_{n}$, and the matrix $\mathcal{H}$ and time $t_{2}$ are such that the second column of unitary matrix $e^{-i\mathcal{H}t_{2}}$ has matrix elements $\left(e^{-i\mathcal{H}t_{2}}\right)_{m2}=\frac{1}{\sqrt{3}}$, we can derive the fermion operators at these two points as
@@
When $\rho_{\text{tot}}$ is a mixed state, we can also define the mutual information between sub-Hilbert space $h_{A}$ of slice $A$ and its complement $h_{{A^{c}}}$:
@@
where $S_{A^{c}}=-\text{tr}(\rho_{A^{c}}\ln\rho_{A^{c}})$ is the entanglement entropy of subsystem $A^{c}$, and $S_{\text{tot}}=-\text{tr}(\rho_{\text{tot}}\ln\rho_{\text{tot}})$ is the entanglement entropy of the entire system. For a pure state $\rho_{\text{tot}}$, one has $S_{\text{tot}}=0$ and $\mathcal{I}=2S_{A}$. We note that however, for mixed states, neither entanglement entropy nor mutual information is a good measure of the entanglement between $h_{A}$ and $h_{{A^{c}}}$. (In fact, for mixed states it is more proper to call $S_{A}$ the von-Neumann entropy, but we retain the name entanglement entropy for convenience, as is adopted in many literature.) The entanglement of mixed states can be for instance characterized by the entanglement negativity [53, 54], which we leave for future studies.
@@ II.2 Path integral formalism
To provide an understanding of the above definition of spacetime slice entanglement entropy in Eq. 17 from a spacetime local perspective, in this subsection we rewrite it in terms of the coherent state path integral formalism. The readers who are solely interested in the physical outcomes of the spacetime slice entanglement entropy in Eq. 17 may skip this subsection.
@@
II.3 Definition for free fermions with pairing
We can generalize our definition of spacetime entanglement entropy for free fermion Bogoliubov-de Gennes (BdG) models with pairing of the form
$H=\sum_{i,j}\left[h_{ij}c^{\dagger}_{\mathbf{r}_{i}}c_{\mathbf{r}_{j}}+\frac{1}{2}\left(\Delta_{ij}c^{\dagger}_{\mathbf{r}_{i}}c^{\dagger}_{\mathbf{r}_{j}}+h.c.\right)\right]\ ,$ (34)
where $\Delta_{ij}=-\Delta_{ji}$ is the pairing amplitude.
In the slice $A$ with $K$ points $(\mathbf{r}_{n},t_{n})$, the annihilation operator $c_{\mathbf{r}_{n}}(t_{n})=e^{iHt_{n}}c_{\mathbf{r}_{n}}e^{-iHt_{n}}$ at point $(\mathbf{r}_{n},t_{n})$ will be a Bogoliubov fermion mode which is the linear combination of both $c_{\mathbf{r}_{m}}$ and $c^{\dagger}_{\mathbf{r}_{m}}$ at time $t=0$. This leads to anti-commutation relations
$\begin{split}&\{c_{\mathbf{r}_{m}}(t_{m}),c_{\mathbf{r}_{n}}^{\dagger}(t_{n})\}=B_{mn}\ ,\\ &\{c_{\mathbf{r}_{m}}(t_{m}),c_{\mathbf{r}_{n}}(t_{n})\}=V_{mn}\ ,\quad(m,n\in A)\ ,\end{split}$ (35)
where by definition $B_{mn}$ is a Hermitian matrix, and $V_{mn}$ is a symmetric matrix (not necessarily real). By defining the Nambu basis $\Phi_{A}=(c_{\mathbf{r}_{1}}(t_{1}),\cdots,c_{\mathbf{r}_{K}}(t_{K}),c_{\mathbf{r}_{1}}^{\dagger}(t_{1}),\cdots,c_{\mathbf{r}_{K}}^{\dagger}(t_{K}))^{T}$ of slice $A$, we can rewrite the above anti-commutation relations as
$\{\Phi_{A},\Phi_{A}^{\dagger}\}=\left(\begin{array}[]{cc}B&V\\ V^{*}&B^{*}\end{array}\right)=R\ ,$ (36)
where $B^{*}$ is the complex conjugation of matrix $B$, and $V^{\dagger}$ is the Hermitian conjugation of matrix $V$. In the absence of pairing $\Delta_{ij}$, one would have $V=0$. By definition, the entire matrix $R$ is non-negative, Hermitian and respects the particle-hole symmetry $PRP^{-1}=R^{*}$, as ensured by $P\Phi_{A}=(\Phi_{A}^{\dagger})^{T}$, where the particle-hole transformation matrix $P=I_{K}\otimes\sigma_{x}$, with $I_{K}$ being the $K\times K$ identity matrix and $\sigma_{x}$ the Pauli-$x$ matrix.
Then, one can show that there exists a transformation matrix $L$ (not unique) such that
$L^{\dagger}L=R^{-1}\ ,\qquad PLP^{-1}=L^{*}\ ,$ (37)
namely, preserves the particle-hole symmetry $P$. This allows us to transform the fermion operators into a new Nambu basis $\Psi_{A}=(d_{1},d_{2},\cdots,d_{K},d_{1}^{\dagger},d_{2}^{\dagger},\cdots,d_{K}^{\dagger})^{T}$ given by
$\Psi_{A}=L\Phi_{A}\ ,$ (38)
where the fermion annihilation and creation operators $d_{j}$, $d_{j}^{\dagger}$ are orthogonal:
$\{d_{i},d_{j}^{\dagger}\}=\delta_{ij}\ ,\qquad\{d_{i},d_{j}\}=\{d_{i}^{\dagger},d_{j}^{\dagger}\}=0\ .$ (39)
A systematical way of obtaining the transformation matrix $L$ satisfying Eq. 37 can be found in the SM [51]. Such a set of $K$ fermion modes $d_{j}$ spans the $2^{K}$ dimensional subHilbert space $h_{A}$ of spacetime slice $A$, similar to the fermion number conserving case.
Once the fermion operators $d_{j}$, $d_{j}^{\dagger}$ in Eq. 37 are obtained, which identifies the degrees of freedom of the subsystem of spacetime slice $A$, we can calculate the entanglement entropy $S_{A}$ following Ref. [52]. For any given quantum state $\rho_{\text{tot}}$ which is a Fock state or generically Gaussian state, one can calculate the entanglement entropy $S_{A}$ inside the spacetime slice $A$ by calculating both the normal and pairing correlation matrices:
$\begin{split}&D_{ij}=\text{tr}\left[\rho_{\text{tot}}d_{i}^{\dagger}d_{j}\right]\ ,\quad F_{ij}=\text{tr}\left[\rho_{\text{tot}}d_{j}^{\dagger}d_{i}^{\dagger}\right]\ ,\end{split}$ (40)
and then calculate the reduced density matrix $\rho_{A}$ by following the standard formula derived in Ref. [52], which we will not repeat here due to its relatively complicated form.
III ENTANGLEMENT ENTROPY IN TIME-DIRECTION SLICE
III.1 Model and definitions
We note that our definition of entanglement entropy in generic spacetime slice above holds for any free fermion models in any spacetime dimensions. For simplicity, hereafter, we investigate the spacetime slice entanglement entropy for the 1D tight-binding free fermion model with fermion number conservation (namely, without pairing).
We consider the following model with $L$ sites and periodic boundary condition:
@@ III.1 Model and definitions
We now consider the case of slice $A$ being a time-direction line segment of total time $t$ on a fixed site $\mathbf{r}_{n}=0$, containing $K$ equally spaced points at time $t_{n}=(n-1)\tau$ with coordinates
$(\mathbf{r}_{n},t_{n})=(0,t_{n})=(0,(n-1)\tau)\ ,\quad(1\leq n\leq K)$ (43)
and $\tau=\frac{t}{K-1}$ is the spacing between adjacent times, as shown in Fig. 1b.
We first study the entanglement entropy $S_{A}$ of thermal states $\rho_{\text{tot}}$ at temperature $T$ and filling $\nu$ in slice $A$, and then study that for non-eigenstates. We set $L\gg K$, for which $L$ can be approximately viewed as infinite (the thermodynamic limit), and $S_{A}$ is independent of $L$.
III.2 Zero temperature
At temperature $T=0$, $\rho_{\text{tot}}$ is a pure state (ground state with a Fermi sea). For fixed total time $t$, as shown in Fig. 3a, within the range $K<t/\tau_{0}$, or equivalently $\tau>\tau_{0}$, $S_{A}$ shows a triangular (trapezoid) shape curve with respect to $K$ for filling $\nu=0.5$ ($\nu\neq 0.5$); while in the range $K>t/\tau_{0}$, or $\tau<\tau_{0}$, $S_{A}$ approaches a stabilized value.
We further examine the total time $t$ dependence of the stabilized value of $S_{A}$ by fixing the time separation $\tau$ and changing the number of points $K=\frac{t}{\tau}+1$. Interestingly, we find $S_{A}$ is approximately linear in $\ln t$, where the slope $\left(\frac{\partial S_{A}}{\partial\ln K}\right)_{\tau}\approx\left(\frac{\partial S_{A}}{\partial\ln t}\right)_{\tau}$ is $\frac{1}{3}$ at $\tau=\tau_{0}$ (Fig. 3c), and quickly saturates to $\frac{1}{6}$ when $\tau<\tau_{0}$ (Fig. 3d). In addition, $S_{A}$ shows an oscillation of period $\text{max}(\frac{1}{\nu},\frac{1}{1-\nu})$ with respect to $K$ [51] at filling $\nu$. These conclusions are insensitive to the energy dispersion [51].
@@ III.2 Zero temperature
Figure 3: (a)-(d) Entanglement entropy $S_{A}$ of model Eq. 41 with $L=500$ in time-direction slice of length $t$ with $K$ points separated by $\tau=t/(K-1)$, for (a) different filling $\nu$ given in the legend at temperature $T=0$ and fixed $1682\pi t/\tau_{0}=$ 168; (b) finite temperature $S_{A}$ with fixed $1682\pi t/\tau_{0}=$ 168, at fillings $\nu=0.5$ (solid) and $\nu=0.3$ (dashed). For each $\nu$, the four curves from low to high have $T/E_{0}=0,0.075,0.15,\infty$, respectively. (c)-(d) $S_{A}$ versus $\ln t$ for different fillings $\nu$ (see legend) at $T=0$ and fixed (c) $\tau=\tau_{0}$ and (d) $\tau=\frac{2}{\pi}\tau_{0}$. (e)-(f) Mutual information $\mathcal{I}$ with $L=200$ at $\nu=0.5$ and temperatures $T/E_{0}$ given in the legends while fixing (e) $1682\pi t/\tau_{0}=$ 168 and (f) $\tau=\frac{5}{2\pi}\tau_{0}$, respectively.
@@
where $n_{F}(E)$ is the zero temperature Fermi-Dirac distribution function at filling $\nu$. The above expressions are generically valid in the infinite system size limit $L\rightarrow\infty$.
III.2.1 The $\tau\gg\tau_{0}$ case: volume law
When $\tau\gg\tau_{0}$, or $K\ll t/\tau_{0}$, one has $|t_{m}-t_{n}|\gg\tau_{0}=2\pi/E_{0}$ for $m\neq n$. This leads to
@@ III.2.1 The $\tau\gg\tau_{0}$ case: volume law
since the off-diagonal elements with $m\neq n$ are suppressed by their fast oscillating integrands proportional to $e^{\mathrm{i}E\left(t_{m}-t_{n}\right)}$. As a result, the $B$ matrix is simply the identity matrix, and thus by Eq. 6 one can choose the matrix $M=I$. This yields $D=M^{*}CM^{T}\approx\nu I$, and by Eq. 17 one has
@@
This fits well with the slope $\left(\frac{\partial S_{A}}{\partial K}\right)_{t}$ at small $K$ in Fig. 3a [51]. Note that this result also holds for any finite temperature, since the finite temperature Fermi-Dirac distribution function $n_{F}(E)$ will not change Eq. 46.
In particular, if one approximates $\Omega(E)$ as uniform within the range $[-\frac{E_{0}}{2},\frac{E_{0}}{2}]$:
$\Omega(E)=E_{0}^{-1}\Theta\left(\frac{E_{0}}{2}-|E|\right)\ ,$ (48)
where $\Theta(x)=1$ ($=0$) when $x>0$ ($x<0$), Eq. 47 becomes exact when $\nu=\frac{p}{q}$ ($p,q\in\mathbb{N}^{+}$ coprime) and $\tau=jq\tau_{0}$ ($j\in\mathbb{N}^{+}$).
Since $t=(K-1)\tau$, Eq. 47 implies that for fixed $\tau\gg\tau_{0}$, the time-direction entanglement entropy has a time volume law (in the infinite system size limit $L\rightarrow\infty$):
$S_{A}\propto t\ ,\qquad(\tau\gg\tau_{0})\,$ (49)
Intuitively, this is because the discrete points $t_{m}$ in the time-direction slice $A$ are too far from each other when $\tau\gg\tau_{0}$, and thus they are nearly uncorrelated (as seen from $C\propto I$) and have almost orthogonal fermion basis (as seen from $B=I$). As a result of lack of correlations, the time-direction entanglement entropy $S_{A}$ is approximately the sum of the entropy of each time point $t_{m}$, and thus obey the volume law when $\tau\gg\tau_{0}$.
We note that $S_{A}$ in this limit is similar to the tensor network temporal entanglement entropy in [49, 50], although the definitions differ significantly [51].
III.2.2 The $\tau=\tau_{0}$ case: duality to the spatial case
When $\tau=\tau_{0}$, or $K=\frac{t}{\tau_{0}}+1$, the time direction entanglement entropy shows a similar behavior as the conventional spatial entanglement entropy. This can be most easily seen by taking the uniform $\Omega(E)$ approximation in Eq. 48, for which one has exactly